40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Dec 5 17:34:30 2009, rob, Update, IOO, frequency noise problem mcf.png
    Reply  Sat Dec 5 22:31:52 2009, rob, Update, IOO, frequency noise problem 
Message ID: 2359     Entry time: Sat Dec 5 22:31:52 2009     In reply to: 2357
Author: rob 
Type: Update 
Category: IOO 
Subject: frequency noise problem 

Quote:

There's a large broadband increase in the MC_F spectrum.  I'm not totally sure it's real--it could be some weird bit-swapping thing.  I've tried soft reboots of c1susvme2 and c1iovme, which haven't helped.  In any case, it seems like this is preventing any locking success today.  Last night it was fine.

 Rebooting c1iovme (by keying off the crate, waiting 30 seconds, and then keying it back on and restarting) has resolved this.  The frequency noise is back to the 'usual' trace.

ELOG V3.1.3-